29. September 2015: Tuesday

LDAP/IdM:

  • LDAP Updates:
    • Two issues since Friday:
      • delete PeopleSoft entry that is not in LDAP
        • It was in the manual intervention needed list, removed it.
      • OSP with different last name.
        • Determined it is the right person.
        • Merged the data into LDAP and created the account.
  • Account Updates:
    • No Issues
  • Development System:
    • manage_data scripts went haywire and wiped out a lot of files yesterday.
    • Working on recovering the files and getting it working again.

Backups:

  • Save Groups:
    • Loads of them failing – Mostly that CoM/IS disk array that always has trouble.
  • Clones:
    • Started the CoM/IS and Disk clones.

Luminis 5:

  • Working on the RollOut and RollBack ldap commands.

Tags:

Leave a Reply

You must be logged in to post a comment.